# Yuxin Feng4, Long Ma2†, Bo Du5, Pheng-Ann Heng1

### v1

@@ 4 Benchmarking and Results
We provide an overview of benchmark construction and evaluation for surgical image restoration (SIR) based on SurgClean dataset. The benchmark includes implementation details, evaluation metrics, training setting, and performance evaluation and analysis of various representative image restoration models, including generic and task-specific image restoration models. All pre-trained weights and evaluation codes of our benchmark are available at https://github.com/PJLallen/Surgical-Image-Restoration.
@@ 4.1 Experimental Setup
Implementation Details. All methods are implemented with PyTorch [22] and trained on two NVIDIA RTX 4090 GPUs with the Adam optimizer. During training, the input image is randomly cropped into 128$\times$128 patches with a batch size of 2 and a total of 200,000 iterations. In addition, we use random crop and horizontal flips as data augmentation. The initial learning rate of all methods is set to default value, and the learning rate is decreased by half every 100k iterations throughout the training process. Additionally, all models are trained on the training set of each sub-scenario and then individually inferred on the corresponding test set on SurgClean. For desmoking, 480 images are employed for training and 120 for testing. For defogging, 256 images are used for training and 64 for testing. For desplashing, the subset divides the training and test samples in an 8:2 ratio.
Evaluation Metrics. In line with general evaluation metrics for image restoration tasks, this benchmark evaluates model performance using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) metrics. Referring to [31], we also adopt non-reference metrics, i.e., Naturalness Image Quality Evaluator (NIQE) and Perceptual Index (PI), to evaluate the restoration results of three sub-degradation scenarios.
Training Setting. As mentioned in Sec. 3.2, our dataset provides unaligned paired reference labels to facilitate the training process of restoration models. To mitigate the impact of unaligned references, all comparison models use pre-trained PWC-Net [26] to estimate the optical flow of unaligned reference $\mathbf{UR}$ so that it aligns with the predicted restored image $\mathbf{P}$:
@@
Finally, $\mathbf{UR}_{warp}$ is spatially aligned with the restored image $\mathbf{P}$. The reconstruction loss $\mathcal{L}_{rec}$ of the restoration model can be expressed as
@@
where $\odot$ is the pixel-wise multiplication, and Mi masks areas with imperfect optical flow to prevent trivial solutions. This allows the model to be trained and evaluated in the same way as the image restoration task with aligned GT. We apply this warping-based training and evaluation strategy to all models in our benchmark to ensure fairness and effectiveness.
Table 3: Quantitative comparison for surgical image restoration with representative generic restoration models on the SurgClean test sets with PSNR $\uparrow$, SSIM $\uparrow$, LPIPS $\downarrow$, NIQE $\downarrow$ and PI $\downarrow$ metrics.
Methods Pubs Desmoking Defogging Desplashing Params
PSNR SSIM LPIPS NIQE PI PSNR SSIM LPIPS NIQE PI PSNR SSIM LPIPS NIQE PI
Restormer [36] CVPR’22 18.470 0.634 0.368 5.373 11.854 18.154 0.619 0.470 6.757 30.084 21.732 0.708 0.302 5.111 25.807 26.13M
FocalNet [8] ICCV’23 18.347 0.655 0.397 5.631 36.676 18.250 0.626 0.474 7.066 41.550 21.740 0.722 0.286 5.557 35.590 3.74M
ConvIR [9] TPAMI’24 18.464 0.649 0.399 5.807 42.004 18.022 0.617 0.471 6.782 33.173 21.745 0.720 0.309 5.476 37.437 14.83M
Fourmer [39] ICML’24 18.208 0.642 0.408 5.418 28.154 17.712 0.610 0.484 6.530 32.162 21.597 0.714 0.300 5.327 30.685 0.68M
MambaIR [12] ECCV’24 18.357 0.649 0.418 6.310 43.049 17.338 0.625 0.524 7.801 48.292 21.908 0.727 0.310 6.411 46.272 4.31M
Histoformer [27] ECCV’24 16.687 0.376 0.640 6.939 13.202 16.183 0.348 0.632 7.393 11.799 20.348 0.637 0.320 5.408 2.348 29.92M
RAMiT [7] CVPR’24 18.348 0.647 0.386 5.841 29.850 18.162 0.622 0.483 7.070 40.173 21.662 0.716 0.305 5.424 34.413 0.30M
AMIR [35] MICCAI’24 18.097 0.637 0.409 6.332 17.141 18.191 0.604 0.521 7.266 48.706 21.759 0.701 0.285 4.962 11.363 23.54M
AST [40] CVPR’24 18.630 0.617 0.405 5.317 9.088 16.572 0.590 0.422 6.461 25.395 21.865 0.720 0.302 4.977 25.109 19.92M
X-Restormer [5] ECCV’24 17.570 0.647 0.398 5.545 37.243 18.148 0.615 0.471 6.972 26.732 21.882 0.727 0.307 5.575 37.966 42.52M
SFHformer [15] ECCV’24 17.965 0.637 0.405 5.688 9.833 18.035 0.613 0.483 7.168 15.285 21.913 0.727 0.296 5.438 34.271 7.67M
MambaIRv2 [13] CVPR’25 18.572 0.638 0.422 6.046 44.250 17.942 0.625 0.519 7.878 53.505 21.995 0.728 0.315 5.761 44.683 0.77M
@@ 4.2 Benchmark Approaches
We conduct a comprehensive evaluation of 22 representative image restoration methods on SurgClean, including 12 generic image restoration models and 10 task-specific image restoration models. Specifically, generic restoration frameworks are built on different infrastructures, including CNN-based ConvIR [9] and FocalNet [8], transformer-based Restormer [36], X-Restormer [5], AST [40], RAMiT [7], and Histoformer [27], state space model-based MambaIR [12] and MambaIRv2 [13], and Fourier-based methods Fourmer [39] and SFHformer [15].
Besides, we employ task-specific frameworks to validate the effectiveness in single degradation scenarios, including MS-CNN [29], Desmoke-LAP [21], IIT-EDC [24], DCP [14], and SelfSVD [31] for desmoking, and C2PNet [38], RIDCP [32], CORUN [10], DEA [6], and CoA [20] for defogging.
@@ 4.3 Benchmark Results
Quantitative Analysis. We conduct a comprehensive quantitative comparison of various representative generic and task-specific image restoration methods on the SurgClean dataset for desmoking, defogging, and desplashing tasks. As shown in Table 3, existing generic restoration methods still face unique challenges when dealing with complex degradation issues in real-world surgical environments. Specifically, AST [40] achieves relatively superior PSNR, NIQE, and PI scores, i.e., 18.630, 5.317, and 9.088, respectively for desmoking, but these results remain insufficient for high-quality reconstruction of smoke degradation during surgery. FocalNet [8] achieves higher PSNR and SSIM values in the defogging task, but lower NIQE and PI values indicate that restored images still fall short of clean images in terms of authenticity and naturalness. For desplashing, MambaIRv2 [13] and SFHformer [15] perform better in terms of PSNR and SSIM metrics, but the results require further improvement to meet the standards for clinical applications. In comparison, the RAMiT framework [7] has only 0.3M model parameters, which provides certain advantages when in deployment on actual surgical equipment. More fine-grained quantitative results can be found in Appendix. D.1.
@@
Fig. 4 illustrates the quantitative performance of task-specific restoration models for desmoking and defogging on the SurgClean test sets. We can see that approaches specialized for desmoking do not yield better restoration performance compared to generic image restoration models, such as MS-CNN [29] with a score of 18.48 compared to 18.63 achieved by AST [40]. For the defogging task, task-specific models also perform slightly worse than generic restoration models, highlighting the uniqueness and complexity of fog degradation in laparoscopic scenes. Overall, although various restoration techniques have developed rapidly in recent years, the specificity of degradation in multiple types of surgical scenarios, e.g., structural diversity and local damage, hinders the generalization of existing restoration models to real surgical environments.
@@
Qualitative Analysis. We provide visual analysis of several restoration methods with favorable quantitative results in typical desmoking, defogging, and desplashing scenarios. As shown on the left side of Fig. 5, most methods exhibit limited effectiveness in removing dense smoke within localized operating areas, and residual smoke still obscures anatomical structures. FocalNet [8] and AST [40] perform better on tissue visibility, but still struggle to recover fine anatomical details. For defogging, restoration models fail to completely suppress fog to provide a clear view, and structural distortion remains in the operating area. Similarly, most approaches are ineffective in eliminating occlusions caused by local splashes. In general, all benchmarking methods show limited performance when addressing various types of degradation in surgical scenarios, highlighting the uniqueness and persistent LSIR challenges. More visualization results can be found in Appendix. D.2.
@@ 5 Beyond Pixel-Level Recovery
In this section, we attempt to look beyond the “pixel-level recovery” perspective from two aspects, including scene analysis between surgical and natural environments and surgical scene parsing.
5.1 Scene Analysis between Surgical and Natural Environments
Although Section 4 explores a direct adaptation strategy by leveraging existing natural image restoration models for surgical scenarios, extensive experiments reveal fundamental incompatibilities. In this part, we try to explore its fundamental cause. As illustrated in Fig. 6 (left), the t-SNE visualization demonstrates significant distributional divergence between natural and surgical scenes (natural haze [17], natural dense haze [1, 2]), indicating inherent domain gaps at the statistical level. Furthermore, we utilized the depth estimation model [34] to perform depth-related scattering degradation analysis (Fig. 6 (right)), highlighting critical imaging distinctions, i.e., laparoscopic fog exhibits localized depth variations with abrupt transitions, whereas natural scenes present gradual depth progressions across expansive environments. These compounded discrepancies, which span both statistical feature spaces and physical imaging degradation characteristics, collectively explain why conventional natural image restoration paradigms struggle in surgical domains.
Figure 6: Scene analysis between surgical and natural environments. t-SNE feature visualization (Left) and depth estimation examples (Middle/Right).
Figure 7: Surgical scene parsing: SAM [16] (Top) and MedSAM [19] (Bottom).
5.2 Surgical Scene Parsing
While visual data quality remains essential from clinical perspectives, the integration of intelligent assistive cues (e.g., scene understanding) proves critical for advancing precision medicine in surgical practice. To systematically evaluate how image restoration quality impacts downstream tasks, we employed the SAM [16] for scene segmentation and its medical variant MedSAM [19] for instrument localization across surgical scenarios. As demonstrated in Fig. 7, although different restoration algorithms have brought improvements in visual quality, their performance in segmentation tasks does not show stable and consistent performance improvement. Especially in surgical scenes with complex structures and severe occlusion, it is still difficult to achieve accurate segmentation of tissue structures and surgical instruments. Therefore, achieving the optimal equilibrium between visual fidelity and downstream task-specific performance emerges as a critical challenge for maximizing the clinical utility of surgical image restoration models. This fundamental trade-off underscores the necessity of semantic-aware restoration frameworks, charting a promising research direction.
6 Discussion and Conclusion
@@
Discussion. The core objective of this study is to construct a multi-type laparoscopic surgical image restoration (LSIR) benchmark dataset for complex surgical scenarios. By establishing a standardized evaluation platform, we aim to drive the development of restoration algorithms to meet clinical needs. Here, we present a deeper analysis and insights into SurgClean benchmark:
@@
1) Suboptimal benchmark performance. Both quantitative and qualitative analyses indicate that existing restoration methods are inadequate for handling the highly complex and variable degradations in surgical environments. Compared to natural image restoration, LSIR poses unique challenges, e.g., artifact blending with the background, localized occlusions, and the entanglement of diseased tissues with degraded areas, which significantly increase the complexity of surgical scenes. We attempted to explore the differences between surgical and natural images from the perspectives of structural depth and semantic understanding to provide insights for future research.
@@
2) Selection of degradation types and data scale. We selected three degradation scenarios (i.e., smoke, fog, and splashing) from 134 real surgical cases that experienced surgeons identified as directly affecting surgical efficiency. The distribution of samples across different degradation types reflects their relative frequency in the full dataset. We plan to expand the dataset with more real-world samples to facilitate further research in the LSIR community.
@@
3) Simultaneous multi-type degradation. While it is theoretically possible for multiple degradation types to occur simultaneously in the same scenario, such combined degradations were rarely observed in our data collection process. This may be due to differing phases and conditions under which the three types of degradation occur. Thus, the first multi-type surgical restoration benchmark only focuses on samples with a single degradation type in a single scene. Future work will explore simultaneous degradations to enhance the LSIR diversity and challenge.
@@
4) Clinical significance. The surgical image restoration task provides surgeons with a clear surgical field of view, thereby significantly reducing operation time, minimizing surgical errors, and improving procedural efficiency. By eliminating unnecessary visual obstructions, surgeons can maintain focus during operations. Beyond enhancing image quality, LSIR also involves reconstructing complex anatomical details, which is critical for precise surgical navigation and shortening postoperative recovery times for patients.

### v2

@@ 4.1 Experimental Setup
Implementation Details. All methods are implemented with PyTorch [paszke2017automatic] and trained on two NVIDIA RTX 4090 GPUs with the Adam optimizer. During training, the input image is randomly cropped into $\times128$ patches with a batch size of 2 and a total of 200,000 iterations. In addition, we use random crop and horizontal flips as data augmentation. The initial learning rate of all methods is set to default value, and the learning rate is decreased by half every 100k iterations throughout the training process. Additionally, all models are trained on the training set of each sub-scenario and then individually inferred on the corresponding test set on SurgClean. For desmoking, 1,700 images are employed for training and 427 for testing. For defogging, 679 images are used for training and 170 for testing. The subset of desplashing divides the training and test sets in an 8:2 ratio.
Table 3: Quantitative comparison for surgical image restoration with representative generic restoration models on the SurgClean test sets. Best and second-best among all methods are highlighted in pink and blue, respectively.
Methods Pubs Desmoking Defogging Desplashing Params
[…]
Restormer [zamir2022restormer] CVPR’22 18.939 0.674 0.362 5.407 29.552 19.036 0.619 0.449 7.166 21.753 21.396 0.718 0.320 5.451 35.970 26.13M
FocalNet [focal2023] ICCV’23 19.244 0.679 0.367 5.933 37.687 19.069 0.628 0.435 5.967 36.892 21.422 0.717 0.320 5.675 38.538 3.74M
ConvIR [cui2024revitalizing] TPAMI’24 19.432 0.678 0.336 5.515 32.126 18.870 0.619 0.426 5.553 28.893 21.331 0.717 0.341 5.661 42.003 14.83M
Fourmer [zhou2023fourmer] ICML’24 19.367 0.675 0.359 5.506 30.973 18.872 0.619 0.489 6.212 27.266 21.422 0.716 0.330 5.653 36.194 0.68M
MambaIR [guo2025mambair] ECCV’24 19.319 0.679 0.355 6.033 43.266 18.872 0.622 0.449 6.331 34.889 21.426 0.722 0.325 6.066 44.614 4.31M
Histoformer [sun2025restoring] ECCV’24 17.666 0.439 0.481 4.700 12.315 16.573 0.346 0.607 6.729 17.305 20.125 0.579 0.355 4.336 4.269 29.92M
RAMiT [choi2024reciprocal] CVPR’24 19.033 0.677 0.371 5.773 38.571 19.019 0.625 0.439 6.146 34.543 21.429 0.718 0.330 5.775 44.203 0.30M
AMIR [yang2024all] MICCAI’24 19.118 0.680 0.369 5.641 40.091 18.871 0.627 0.440 6.075 32.879 21.292 0.717 0.346 5.441 38.380 23.54M
AST [zhou2024AST] CVPR’24 19.182 0.635 0.382 5.085 9.012 17.046 0.606 0.448 6.182 24.973 22.046 0.731 0.288 4.752 24.906 19.92M
X-Restormer [chen2024xrestormer] ECCV’24 18.034 0.659 0.372 5.201 36.810 18.603 0.628 0.446 6.605 26.084 22.317 0.735 0.294 5.331 37.425 42.52M
SFHformer [jiang2024fast] ECCV’24 19.307 0.668 0.351 5.048 15.992 18.794 0.612 0.432 5.893 18.287 21.231 0.696 0.278 4.636 17.008 7.67M
MambaIRv2 [guo2024mambairv2] CVPR’25 19.391 0.680 0.367 5.758 39.578 18.930 0.624 0.436 6.263 35.709 21.398 0.718 0.325 5.881 40.824 0.77M
Figure 6: The effect of our warping operations, which yields better alignment between paired GTs and input images.
Evaluation Metrics. In line with general evaluation metrics for image restoration, we evaluate model performance using Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS) metrics. Referring to [wu2025self], we also adopt non-reference metrics, i.e., Naturalness Image Quality Evaluator (NIQE) and Perceptual Index (PI), to evaluate the restoration results of three sub-degradation scenarios.
Training Setting. As mentioned in Sec. 3.2, our dataset provides unaligned paired reference labels to facilitate the training process of restoration models. To mitigate the impact of unaligned references, all comparison models use pre-trained PWC-Net [sun2018pwc] to estimate the optical flow of unaligned reference $\mathbf{UR}$ so that it aligns with the predicted restored image $\mathbf{P}$:
@@
The visualization of the warping operations is shown in Fig. 6. Finally, $\mathbf{UR}_{warp}$ is spatially aligned with the restored image $\mathbf{P}$. The reconstruction loss can be expressed as
@@
where $\odot$ is the pixel-wise multiplication, and Mi masks areas with imperfect optical flow to prevent trivial solutions. This allows the model to be trained and evaluated in the same way as the image restoration task with aligned GT. We apply this warping-based training and evaluation strategy to all models in our benchmark to ensure fairness.
Figure 8: Cross-dataset validation for desmoking between DesmokeData [xia2024new] and SurgClean datasets (Training $\rightarrowTesting dataset$).
@@ 4.2 Benchmark Approaches
We conduct a comprehensive evaluation of 22 representative methods on SurgClean, including 12 generic and 10 task-specific models. Specifically, generic frameworks are built on different infrastructures, including CNN-based ConvIR [cui2024revitalizing] and FocalNet [focal2023], transformer-based Restormer [zamir2022restormer], X-Restormer [chen2024xrestormer], AST [zhou2024AST], RAMiT [choi2024reciprocal], and Histoformer [sun2025restoring], state space model-based MambaIR [guo2025mambair] and MambaIRv2 [guo2024mambairv2], and Fourier-based methods Fourmer [zhou2023fourmer] and SFHformer [jiang2024fast]. Besides, we employ task-specific frameworks to validate the effectiveness in single degradation scenarios, including MS-CNN [wang2019multiscale], Desmoke-LAP [pan2022desmoke], IIT-EDC [salazar2020desmoking], DCP [he2010single], and SelfSVD [wu2025self] for desmoking, and C2PNet [zheng2023curricular], RIDCP [wu2023ridcp], CORUN [fang2024real], DEA [chen2024dea], and CoA [ma2025coa] for defogging.
@@ 4.3 Benchmark Results
Quantitative Analysis. We conduct a comprehensive comparison of representative generic and task-specific image restoration methods on SurgClean for three tasks.
@@
As shown in Table 3, existing generic methods still face unique challenges when dealing with complex degradation issues in real-world surgical environments. Specifically, ConvIR [cui2024revitalizing] achieves relatively superior PSNR and LPIPS scores, i.e., 19.432 and 0.336, for desmoking, but these results remain insufficient for high-quality reconstruction of smoke degradation during surgery. FocalNet [focal2023] achieves higher PSNR and SSIM values in defogging, but lower NIQE and PI values indicate that restored images still fall short of clean images in terms of authenticity and naturalness. For desplashing, X-Restormer [chen2024xrestormer] and AST [zhou2024AST] perform better in terms of PSNR and SSIM metrics, but the results require further improvement to meet the standards for clinical applications. In comparison, RAMiT [choi2024reciprocal] has only 0.3M parameters, which provides advantages when deployed on edge equipment. More fine-grained quantitative results can refer to Supp. D.1.
Table 4: Quantitative results on downstream tasks. All metrics were evaluated on the defogging samples of the SurgClean dataset.
Method Scene Analysis: Depth Estimation Scene Parsing: Semantic (Left), Instrument (Right)
[…]
Restormer [zamir2022restormer] 0.3666 1.0348 2.3727 0.4855 0.7344 0.8569 0.4161 0.5869 0.7645 0.8799
FocalNet [focal2023] 0.3359 0.9202 2.2360 0.5234 0.7615 0.8740 0.4112 0.5830 0.7526 0.8716
ConvIR [cui2024revitalizing] 0.3302 0.8885 2.2030 0.5225 0.7680 0.8780 0.4194 0.5877 0.7616 0.8797
Fourmer [zhou2023fourmer] 0.3536 0.9653 2.3477 0.5024 0.7436 0.8628 0.3929 0.5666 0.7593 0.8748
Histoformer [sun2025restoring] 0.4136 1.2328 2.5436 0.4362 0.6935 0.8294 0.2955 0.5284 0.7334 0.8817
RAMiT [choi2024reciprocal] 0.3326 0.8978 2.2273 0.5290 0.7624 0.8764 0.4074 0.5790 0.7646 0.8792
AMIR [yang2024all] 0.3428 0.9370 2.2745 0.5167 0.7527 0.8716 0.3811 0.5652 0.7690 0.8844
AST [zhou2024AST] 0.4096 1.2583 2.5460 0.4368 0.6894 0.8285 0.2600 0.5450 0.7400 0.8950
MambaIR [guo2025mambair] 0.3400 0.9127 2.2698 0.5103 0.7503 0.8684 0.3876 0.5651 0.7642 0.8803
X-Restormer [chen2024xrestormer] 0.3338 0.9876 2.2650 0.5110 0.7502 0.8701 0.3350 0.5250 0.7550 0.8840
SFHformer [jiang2024fast] 0.3400 0.9158 2.2550 0.5137 0.7535 0.8688 0.3915 0.5681 0.7573 0.8737
MambaIRv2 [guo2024mambairv2] 0.3348 0.9103 2.2384 0.5244 0.7611 0.8723 0.4204 0.5902 0.7551 0.8745
Figure 9: Scene analysis between natural and surgical environments. The t-SNE visualization (Left) and depth examples (Middle/Right).
Figure 10: Surgical scene parsing: SAM [kirillov2023segment] (Top) and MedSAM [ma2024segment] (Bottom).
@@
Fig. 5 illustrates the quantitative performance of task-specific restoration models for desmoking and defogging on the SurgClean test sets. We can see that approaches specialized for desmoking do not yield better restoration performance compared to generic image restoration models, such as MS-CNN [wang2019multiscale] with a score of 19.02 compared to 19.43 achieved by ConvIR [cui2024revitalizing]. For the defogging task, task-specific models also perform slightly worse than generic restoration models, highlighting the uniqueness and complexity of fog degradation in endoscopic scenes. Overall, although various restoration techniques have developed rapidly in recent years, the specificity of degradation in multiple types of surgical scenarios, e.g., structural diversity and local damage, hinders the generalization of existing restoration models to real surgical environments.
@@
Qualitative Analysis. As shown in Fig. 7, most methods exhibit limited effectiveness in removing dense smoke within localized operating areas, and residual smoke still obscures anatomical structures. FocalNet [focal2023] and AST [zhou2024AST] perform better on tissue visibility, but still struggle to recover fine anatomical details. For defogging, restoration models fail to completely suppress fog to provide a clear view, and structural distortion remains in the operating area. Similarly, most approaches are ineffective in eliminating occlusions caused by local splashes. In general, all benchmarking methods show limited performance when addressing various types of degradation in surgical scenarios, highlighting the uniqueness and persistent ESIR challenges. More visualization results can be found in Supp. D.2.
4.4 Cross-dataset Validation
[…]
@@ 5 Beyond Pixel-Level Recovery
In this section, we attempt to look beyond the “pixel-level recovery” perspective from two aspects for surgical restoration, including scene analysis and scene parsing.
5.1 Scene Analysis
Table 4 (left) summarizes depth estimation results. Although FocalNet [focal2023] and Histoformer [sun2025restoring] achieve top defogging scores in Table 3, they fail to consistently lead in depth estimation. Strong restoration performance therefore does not guarantee superior depth estimation quality. To explain this mismatch, we examine the underlying domain shift. As shown in Fig. 9 (left), t-SNE reveals a clear distribution gap between surgical scenes and natural haze datasets [li2018benchmarking, ancuti2019ntire, ancuti2020nh]. Using a depth estimator [yang2024depth], our scattering analysis in Fig. 9 (right) further shows that endoscopic fog exhibits localized depth changes with abrupt transitions, whereas natural scenes display gradual, large-scale depth variations. These differences at both the statistical and imaging levels clarify why methods designed for natural image transfer poorly to surgical imagery.
5.2 Scene Parsing
We evaluated scene parsing with SAM for scene segmentation and MedSAM for instrument localization across surgical scenarios. As shown in Table 4 (right), MambaIRv2 achieves better performance on two of four parsing metrics, yet its restoration scores in Table 3 are relatively unremarkable, suggesting that higher restoration scores do not reliably yield stronger parsing. Consistent with this observation, Fig. 10 shows that although restoration enhances visual quality, segmentation improvements remain unstable, particularly in structurally complex scenes. Balancing visual fidelity and task performance is thus crucial for clinical utility, motivating semantic-aware restoration frameworks.
6 Discussion
@@
The core objective of this study is to construct a multi-type endoscopic surgical image restoration (ESIR) benchmark for complex surgical scenarios. By establishing a standardized evaluation platform, we aim to advance restoration algorithms toward clinical applicability. Below are further analyses and insights into the SurgClean benchmark:
@@
1) Suboptimal benchmark performance. Both quantitative and qualitative analyses indicate that existing restoration methods are inadequate for handling the highly complex and variable degradations in surgical environments. Compared to natural image restoration, ESIR poses unique challenges, e.g., artifact blending with the background, localized occlusions, and the entanglement of diseased tissues with degraded areas, which significantly increase the complexity of surgical scenes. We attempted to explore the differences between surgical and natural images from the perspectives of structural depth and semantic understanding to provide insights for future research.
@@
2) Selection of degradation types and data scale. We selected three degradation scenarios (i.e., smoke, fog, and splashing) from 414 real surgical cases that experienced surgeons identified as directly affecting surgical efficiency. The distribution of samples across different degradation types reflects their relative frequency in the full dataset. We plan to continuously expand data with more real-world samples to facilitate further research in the ESIR community.
@@
3) Simultaneous multi-type degradation. While it is theoretically possible for multiple degradation types to occur simultaneously in the same scenario, such combined degradations were rarely observed in our data collection process. This may be due to the differing phases and conditions under which the three types of degradation occur. Thus, the first multi-type surgical restoration benchmark only focuses on samples with a single degradation type in a single scene. Future work will explore simultaneous degradations to enhance the ESIR diversity and challenge.
@@
4) Clinical significance. The surgical image restoration task provides surgeons with a clear surgical field of view, thereby significantly reducing operation time, minimizing surgical errors, and improving procedural efficiency. By eliminating unnecessary visual obstructions, surgeons can maintain focus during operations. Beyond enhancing image quality, ESIR also involves reconstructing complex anatomical details, which is critical for precise surgical navigation and shortening postoperative recovery times for patients.